Find two consecutive integers such that the sum of the first and 3 times the second is 55.
Answer by orca(409)

Let the two consecutive numbers be n, n+ 1.
The sum of the first and 3 times the second can be expressed as
n + 3(n + 1)
Setting it equal to 55, we have
n + 3(n + 1) = 55
Solving for x, we have
n + 3n + 3 = 55
4n + 3 = 55
4n = 52
n = 13
So the two consecutive numbers are 13 and 14.